OpenStudy (anonymous):

Lawson opened a savings account 24 years ago with a deposit of $4,169.06. The account has an interest rate of 4.3% compounded quarterly. How much interest has Lawson earned? Answer $11,636.92 $7,467.86 $11,451.48 $4,348.33

OpenStudy (mathteacher1729):

This is a very very useful page for computing such numbers. http://www.wfu.edu/~palmitar/Law&Valuation/chapter%201/1-2-2.htm Also -- these computations are SUPER important! Understanding your bank's different savings accounts are very important for managing your money.

OpenStudy (tkhunny):

Only one payment. This should not be tripping you up. They only get harder from here. Find a way to be systematic with Basic Principles Annual Interest: 0.043 Quarterly Interest: 0.043/4 = 0.01075 Quarterly Accumulation Factor: 1.01075 Yearly Accumulation Factor: 1.01075^4 = 1.04369835754219140625 Round howver you think it appropriate. 24 Year Accumulation Factor: 1.043698^24 = 2.791235 One more thing to do. I haven't used the initial deposit, yet.

OpenStudy (anonymous):

$7,467.86?